# RoDE: Linear Rectified Mixture of Diverse Experts for Food Large Multi-Modal Models

### v1

@@ Introduction
We conduct an ablation study to evaluate the impact of different LoRA rank configurations. The results, as presented in Table 6, span multiple expert configurations, including: 1) Var1 - Four LoRAs, each with a uniform rank of 5. 2) Var2 - Four LoRAs, each with a uniform rank of 8. 3) Var3 - A heterogeneous set of four LoRAs with ranks of [2, 4, 6, 7]. 4) Var4 - Another diverse set of four LoRAs with a distinct rank composition from Var3, specifically [2, 4, 8, 16]. The analysis of Var1 and Var2 suggests that configurations with higher-ranked LoRA experts are more effective. However, despite having the same number of trainable parameters as Var1, Var3 demonstrates superior performance, indicating that a mix of higher and lower-ranked LoRA experts can also be effective. Moreover, Var3 shows comparable performance to Var2, with only a slight reduction in SacreBLEU score (by 0.08) and Rouge-L score (by 0.36), while having 32.5% fewer trainable parameters. This finding supports the notion that not all LoRA experts require high capacity to contribute effectively.
@@
Training Parameter Efficiency
@@
Conclusion

### v3

@@ V-C3 Rank configuration of LoRA
We conduct an ablation study to evaluate the impact of different LoRA rank configurations. The results, as presented in Table VII, span multiple expert configurations, including: 1) Var1 - Four LoRAs, each with a uniform rank of 5. 2) Var2 - Four LoRAs, each with a uniform rank of 8. 3) Var3 - A heterogeneous set of four LoRAs with ranks of [2, 4, 6, 7]. 4) Var4 - Another diverse set of four LoRAs with a distinct rank composition from Var3, specifically [2, 4, 8, 16]. The analysis of Var1 and Var2 suggests that configurations with higher-ranked LoRA experts are more effective. However, despite having the same number of trainable parameters as Var1, Var3 demonstrates superior performance, indicating that a mix of higher and lower-ranked LoRA experts can also be effective. Moreover, Var3 shows comparable performance to Var2, with only a slight reduction in SacreBLEU score (by 0.08) and Rouge-L score (by 0.36), while having 32.5% fewer trainable parameters. This finding supports the notion that not all LoRA experts require high capacity to contribute effectively.
@@
V-D Additional Analysis
V-D1 Training Parameter Efficiency
@@
V-D2 Empirical Analysis
We present some examples of ingredient recognition on FoodLMM and RoDE, as depicted in Figure 11. In these examples, RoDE demonstrates higher accuracy in ingredient recognition. However, it still incorrectly identifies some ingredients, and tends to predict commonly used but ambiguous ingredients.
Figure 11: Comparison between FoodLMM and RoDE on ingredient recognition.
V-D3 Computational Overhead
RoDE employs the rectified linear union, a strategy tailored for sparse expert activation, which is efficient for both training and inference. The average training time for a single batch and the average inference time for recipe generation on the UniFood dataset are demonstrated in Table VIII. Remarkably, while RoDE exhibits superior performance compared to the standard MoE design (as shown in Table  III and Table  VI), it manages to maintain a time consumption that is comparable to that of the standard MoE design.
TABLE VIII: Training and inference computation cost. The training procedure utilizes a batch size of 4 and is conducted on four RTX4090 GPUs. While the inference process is carried out on a single RTX4090 GPU.
FoodLMM FoodLMM+MoE RoDE
Training per batch 21.87s 23.68s 23.72s
Inference on recipe generation 19.21s 20.41s 20.31s
VI Conclusion